# Scaling Efficient Masked Autoencoder Learning on Large Remote Sensing Dataset

### v1

@@ 4.1 Masked Autoencoders Preliminaries
1) Masking. Similar to supervised training of a standard ViT, MAE divides the image into regular, non-overlapping patches. It then samples a subset of these patches and masks the remaining ones. Typically, the masking ratio is 75%, meaning only 25% of the patches are input to the encoder. This random sampling follows a uniform distribution according to the masking ratio. 2) MAE Encoder. The encoder is a standard ViT applied only to the visible, unmasked patches. It linearly projects the patches, adds positional embeddings, and processes them through a series of transformer blocks. By operating on a smaller subset of patches, the encoder enables training of large models with reduced computational and memory requirements. 3) MAE Decoder.The encoded tokens and masked tokens are fed into the decoder, which comprises transformer blocks with self-attention layers. The masked tokens are shared, learnable tensors enhanced with positional embeddings. The decoder, utilized only during pre-training, generates the output predictions for those masked tokens. 4) Reconstruction Target. MAE predicts the pixel values for each masked patch, with each element in the decoder output representing a patch’s pixel value vector. The loss function computes the mean squared error (MSE) between the reconstructed targets and original patches.
@@
RS optical images typically contain many redundant background pixels, which aligns well with MAE’s masking strategy. In MAE, the encoder processes only 25% of the patches, significantly reducing computational load by avoiding processing the entire image. Building on MAE, we aim to leverage the redundancy in RS images to accelerate training. Specifically, we address two questions: 1) Is it necessary to reconstruct all the masked patches given the redundancy in RS images? 2) Can the visible patches input to the MAE encoder be further compressed to enhance acceleration?
@@
4.2 Partial Reconstruction
Table 2: Results of different partial reconstruction strategies. Reconstruction Ratio=25%. Top-1 classification accuracy on RESISC-45 [77] is reported.
Method Selection Data Throu. Acc.
/Min.
Partial Reco. False 17k 89.2
Partial Reco. Random 25k 88.8
Partial Reco. ATS [78] 20k 85.7
Partial Reco. HOG 25k 89.2
For question 1, previous research [45] has shown that for general images, when MAE reconstructs 75% of the patches to calculate the loss, a specially designed decoder doesn’t need to fully reconstruct all remaining patches. In fact, reconstructing just 50% or even 25% of the patches can achieve similar performance and speed up training. However, for RS images, if we randomly sample patches and remove most for reconstruction, the reconstructed patches might not be semantically rich ones. As shown in Table 2, using only a random subset for reconstruction degrades performance.
To address this issue, we propose selecting semantically rich patches for reconstruction instead of random selection. Specifically, given an input image $x\in\mathbb{R}^{H\times W\times C}$, it is reshaped into $N=(H\times W)/p^{2}$ non-overlapping patches $x^{p}\in\mathbb{R}^{N\times(p^{2}C)}$, where $p$ is the patch size, $(H,W)$ is the size of the input image, and $C$ is the number of channels. These patches $\{x_{i}^{p}\}_{i=1}^{N}$ are then linearly mapped to patch embeddings. To retain positional information, positional embeddings are added to the patches. We select a portion of the patches to input to the encoder based on the masking ratio $m\in[0,1]$ ($m=85\%$ by default), as detailed in Sec. 4.1. The remaining patches serve as reconstruction targets for the decoder. Unlike MAE’s masking ratio, we introduce a new reconstruction ratio $r$, the proportion of pixels to be reconstructed, denoted as $r\in[0,m]$ ($r=25\%$ by default). We compute the HOG features $HOG(\cdot)$ of the remaining patches and select those with the high HOG feature values according to the reconstruction ratio $r$, rather than using all patches. The process can be formulated as:
@@ 4.2 Partial Reconstruction
where $token_{R}$ denotes the selected mask tokens for reconstruction and $\text{top}_{n}(\cdot)$ denotes the index set of the selected top $n$ tokens. The decoder uses a lightweight design based on cross-attention following CrossMAE [45]. Experimental results in Table 2 show that this partial reconstruction strategy significantly increases the training throughput without affecting the learned representations.
4.3 Progressive Semantic Token Selection
For the second question, we initially tried a naive approach by increasing the masking ratio to 85%, meaning only 15% of the patches in each RS image were input to the encoder while keeping a 25% reconstruction ratio as proposed in Sec. 4.2. However, during training, this setup often led to issues like gradient explosions or loss divergence, as shown in Fig. 3. To illustrate, we continued MAE’s training despite these gradient explosions. The figure shows that using only 40% (=15%+25%) of the patches for encoding and reconstruction frequently caused unstable training due to gradient explosions. We speculate that, although RS images have background redundancy, using such a small portion of patches significantly challenges MAE training.
Figure 3: Left: Baseline training loss. Using only 40% of masked patches for reconstruction often leads to gradient explosions. Right: Training loss using PSTS.
How can we achieve acceleration at a high masking ratio (e.g., 85%) while ensuring MAE completes the pre-training task for RS images? Some studies suggest that machine learning algorithms often benefit from starting with simpler tasks and gradually increasing the difficulty, a strategy known as curriculum learning [79, 80, 81]. This approach involves progressively introducing more challenging examples during training. While curriculum learning typically focuses on whole image samples, self-supervised training, which masks large portions of image pixels, emphasizes patch-level selection. Therefore, in self-supervised training, it’s essential to distinguish between “easy" and “hard" patches.
@@ 4.3 Progressive Semantic Token Selection
For initialization, we employ a HOG selection strategy to choose the initial patch set from $S^{N}=\left\{x_{i}^{p}\right\}_{i=1}^{N}$, with a proportion $s\in[0,(1-m)/2]$. We define the initial set of selected token as:
[…]
We select $\lfloor\ s\times N\rfloor$ tokens with the maximum HOG feature values (i.e., ${top}_{\lfloor\ s\times N\rfloor}$) from the original token set to form the initial token set. This simple yet effective strategy ensures that semantically rich tokens are selected.
Following token initialization, we aim to incrementally increase the number of tokens to facilitate a training progression from easy to difficult, while adhering to the final masking ratio for selected tokens. This process is outlined in Algorithm 1. Given the initial token sets’ high HOG feature values, nearby tokens selected by PSTS also exhibit high HOG feature values. Recalling our partial reconstruction approach in Section 4.2, we filter the remaining unselected token set to include those with high HOG values as reconstruction targets. Consequently, when selecting nearby tokens, the tokens input for encoding and those chosen for reconstruction are closer, thereby simplifying the training process. Conversely, if a token’s feature distribution significantly deviates from others in the set, indicating it is farther away, we consider such patches more challenging to learn.
Specifically, we select tokens from $S^{U}$ based on $S^{I}$. First, we use $\mathbb{S^{I}}\in\mathbb{R}^{|S^{K}|\times d}$ and $\mathbb{S^{U}}\in\mathbb{R}^{|S^{U}|\times d}$ to denote the matrix representation of the initial token set $S^{I}$ and the unselected token set $S^{U}$, where ||$\cdot$|| ⋅ | represents the number of tokens and $d$ the feature dimension after the embedding layer. We use Cosine Distance to measure the distance between the tokens in these two sets:
$\mathcal{D}(S^{U},S^{I})=\mathbf{1}-cos\left\langle S^{U},S^{I}\right\rangle=\mathbf{1}-\frac{\mathbb{S^{U}}\mathbb{S^{I}}^{\mathrm{T}}}{\left|\kern-1.07639pt\left|\mathbb{S^{U}}\right|\kern-1.07639pt\right|\cdot\left|\kern-1.07639pt\left|\mathbb{S^{I}}\right|\kern-1.07639pt\right|},$ (3)
where $\mathbf{1}$ is an all-one matrix. $\mathcal{D}(S^{U},S^{I})\in\mathbb{R}^{|S^{U}|\times|S^{I}|}$ represents the pairwise distances between tokens in $S^{U}$ and $S^{I}$. Next, we define the distance between the tokens in $S^{U}$ to the initial token set $S^{K}$ based on the selection criteria in each training stage as follows:
[…]
where $i\in\{1,\cdots,|S^{U}|\}$, $j\in\{1,\cdots,|S^{I}|\}$, and $\zeta$ represents the training stage depending on the number of epochs. Finally, we sample $\lfloor N\times(1-m-s)\rfloor$ tokens from ${S^{U}}$ and add them with ${S^{I}}$ to form ${S^{K}}$, which can be formulated as follows:
[…]
${S^{K}}={S^{I}}\cup\mit{S}^{*},\quad{S^{U}}={S^{U}}\setminus\mit{S}^{*},$ (6)
where $\mit{S}^{*}$ represents the selected token from ${S^{U}}$. The operations in Eq. (5) and Eq. (6) are performed in each training epoch. This process is summarized in Algorithm 1.
Algorithm 1 Progressive Semantic Token Selection
1:Number of training epochs $T$ and total training stages $N_{g}$ (i.e., $T/N_{g}$ epochs for each stage), masking rate $m$, input dataset $\mathcal{X}$
2:Obtain the selected tokens set ${S^{K}}$ and update ${S^{U}}$ in each epoch
3:for $t\leftarrow 1$ to $T$ do
4:     Sample data sample from $\mathcal{X}$, feed-forwarded through the embedding to obtain the output token set ${S}_{N}$
5:     $s\leftarrow\frac{1}{2}(1-m)$
6:     Obtain ${S}^{I}$ via Eq. (2) and initialize ${S^{U}}$
7:     Obtain the current training stage $\zeta=\lceil N_{g}*t/T\rceil$
[…]
9:     Obtain $S^{*}$ and update $S^{K},S^{U}$ via Eq. (5) and Eq. (6)
10:end for
@@ 5 Experiment
In this section, we first discuss the pre-training process using the RS-4M dataset, comparing the training speed and performance of our proposed SelectiveMAE with existing MAE acceleration methods. We then demonstrate SelectiveMAE’s superior performance across various downstream tasks compared to different RSFMs. Finally, we explore design choices such as the reconstruction ratio, mask ratio, and decoder depth through extensive ablation experiments, detailed in the Appendix.
5.1 Pre-training
Pre-training Setup. Our pre-training experiment setup largely follows MAE [10]. Unlike other MIM methods in the RS field, such as SatMAE [30] and ScaleMAE [33], we train both the commonly used ViT-B and the larger ViT-L backbones [39]. We adjust the input image to a size of $224\times 224$, with a patch size of 16, using the AdamW optimizer and a cosine learning rate scheduler. The initial learning rate is set to 1.5e-4, batch size to 1024, masking ratio to 85%, and reconstruction ratio to 25%. For patch selection during the 800 epochs, the adjustment thresholds are set at 300 epochs, with the last 200 epochs using random distance to enhance robustness.
Table 3: Top-1 classification accuracy on RESISC-45 [77]. “-” indicates unavailable accuracy data. Training throughput was measured using 8 NVIDIA 40G A100 GPUs.
Model Publication Backbone Params Data Throughput/Min. Acccuracy
MAE [10] CVPR’22 ViT-B 86M 17k 93.45
CrossMAE [45] Arxiv’24 ViT-B 86M 22k 92.82
SatMAE [30] NIPS’22 ViT-B 86M 15k -
ScaleMAE [33] ICCV’23 ViT-B 86M 16k -
SelectiveMAE - ViT-B 86M 37k ($\times2$) 94.58 ($+1.13$)
MAE [10] CVPR’22 ViT-L 307M 15k 93.80
SatMAE [30] NIPS’22 ViT-L 307M 13k 94.10
ScaleMAE [33] ICCV’23 ViT-L 307M 13k 95.04
SelectiveMAE - ViT-L 307M 41k (2.7×2.7$\times$2.7 ×) 95.77 ($+1.97$)
Pre-training Results.
We pre-trained ViT-B and ViT-L on the RS-4M dataset, evaluating the image data throughput per minute for SelectiveMAE and the baseline (MAE) [10]. We also assessed the training speed of the general domain acceleration method CrossMAE [45], as well as the remote sensing methods SatMAE [30] and ScaleMAE [33]. Table 3 illustrates our method’s more than doubled acceleration for both base and large ViT models compared to the vanilla MAE baseline. It also outperforms the general domain acceleration method CrossMAE significantly. Additionally, we reported the top-1 classification accuracy on the same dataset, with both MAE and CrossMAE trained on MillionAID [20]. Table 3 clearly demonstrates our method’s superior speed and performance.
The significant acceleration effect of SelectiveMAE stems from utilizing only 40% of the image patches (15% for the encoder and 25% for the decoder). Despite this reduced utilization, it achieves superior performance compared to other acceleration methods [45] due to the proposed novel token masking and reconstruction strategy, tailored for the remote sensing images. This unique approach ensures that SelectiveMAE delivers both speed and high performance, as validated in Table 4.
5.2 Fine-tuning on Downstream Tasks

### v3

@@ 3.2.1 Masked Autoencoders Preliminaries
1) Masking. Similar to supervised training of a standard ViT, MAE divides the image into regular, non-overlapping patches. It then samples a subset of these patches and masks the remaining ones. Typically, the masking ratio is 75%, meaning only 25% of the patches are input to the encoder. This random sampling follows a uniform distribution according to the masking ratio. 2) MAE Encoder. The encoder is a standard ViT applied only to the visible, unmasked patches. It linearly projects the patches, adds positional embeddings, and processes them through a series of transformer blocks. By operating on a smaller subset of patches, the encoder enables the training of large models with reduced computational and memory requirements. 3) MAE Decoder.The encoded tokens and masked tokens are fed into the decoder, which comprises transformer blocks with self-attention layers. The masked tokens are shared, learnable tensors enhanced with positional embeddings. The decoder, utilized only during pre-training, generates the output predictions for those masked tokens. 4) Reconstruction Target. MAE predicts the pixel values for each masked patch, with each element in the decoder output representing a patch’s pixel value vector. The loss function computes the mean squared error (MSE) between the reconstructed targets and original patches.
@@
RS optical images typically contain many redundant background pixels, which aligns well with MAE’s masking strategy. In MAE, the encoder processes only 25% of the patches, reducing computational load by avoiding processing the entire image. Building on MAE, we aim to leverage the redundancy in RS images to accelerate training. Specifically, we address two questions: 1) Is it necessary to reconstruct all the masked patches given the redundancy in RS images? 2) Can the visible patches input to the MAE encoder be further compressed to enhance acceleration?
@@
3.2.2 Partial Reconstruction
For question 1, previous research [45] has shown that for general images, when MAE reconstructs 75% of the patches to calculate the loss, a specially designed decoder doesn’t need to fully reconstruct all remaining patches. In fact, reconstructing just 50% or even 25% of the patches can achieve similar performance and speed up training. However, for RS images, if we randomly sample patches and remove most for reconstruction, the reconstructed patches might not be semantically rich ones. Using only a random subset for reconstruction even degrades performance.
To address this issue, we propose selecting semantically rich patches for reconstruction instead of random selection. Specifically, given an input image $x\in\mathbb{R}^{H\times W\times C}$, it is reshaped into $N=(H\times W)/p^{2}$ non-overlapping patches $x^{p}\in\mathbb{R}^{N\times(p^{2}C)}$, where $p$ is the patch size, $(H,W)$ is the size of the input image, and $C$ is the number of channels. These patches $\{x_{i}^{p}\}_{i=1}^{N}$ are then linearly mapped to patch embeddings. To retain positional information, positional embeddings are added to the patches. We select a portion of the patches to input to the encoder based on the masking ratio $m\in[0,1]$ ($m=85\%$ by default), as detailed in Sec. 3.2.1. The remaining patches serve as reconstruction targets for the decoder. Unlike MAE’s masking ratio, we introduce a new reconstruction ratio $r$, the proportion of pixels to be reconstructed, denoted as $r\in[0,m]$ ($r=25\%$ by default). We compute the HOG features $HOG(\cdot)$ of the remaining patches and select those with the high HOG feature values according to the reconstruction ratio $r$, rather than using all patches. The process can be formulated as:
@@ 3.2.2 Partial Reconstruction
where $token_{R}$ denotes the selected mask tokens for reconstruction and $\text{top}_{n}(\cdot)$ denotes the index set of the selected top $n$ tokens. The decoder uses a lightweight design based on cross-attention following CrossMAE [45]. Experimental results in the Appendix show that this partial reconstruction strategy significantly increases the training throughput without affecting the learned representations.
3.2.3 Progressive Semantic Token Selection
For the second question, we initially tried a naive approach by increasing the masking ratio to 85%, meaning only 15% of the patches in each RS image were input to the encoder while keeping a 25% reconstruction ratio as proposed in Sec. 3.2.2. However, during training, this setup often led to issues like gradient explosions or loss divergence (see the Appendix). We speculate that, although RS images have background redundancy, using such a small portion of patches challenges MAE training. How can we achieve acceleration at a high masking ratio while ensuring MAE convergence? Curriculum learning [82, 83, 84] may offer a solution. However, it typically focuses on selecting whole image samples, whereas MIM pre-training masks large portions of image pixels, emphasizing patch-level selection. Therefore, it’s essential to distinguish between “easy” and “hard” patches in MIM pre-training.
@@ 3.2.3 Progressive Semantic Token Selection
For initialization, we employ a HOG selection strategy to choose the initial patch set. This simple yet effective strategy ensures that semantically rich tokens are selected. Following token initialization, we aim to incrementally increase the number of tokens to facilitate a training progression from easy to difficult, while adhering to the final masking ratio for selected tokens. This process is outlined in the Appendix. Given the initial token sets’ high HOG feature values, nearby tokens selected by PSTS also exhibit high HOG feature values. Recalling our partial reconstruction approach in Sec. 3.2.2, we filter the remaining unselected token set to include those with high HOG values as reconstruction targets. Consequently, when selecting nearby tokens, the tokens input for encoding and those chosen for reconstruction are closer, thereby simplifying the training process. Conversely, if a token’s feature distribution deviates from others in the set, indicating it is farther away, we consider such patches more challenging to learn. Details of the process are presented in the Appendix.
@@ 4 Experiment
In this section, we first discuss the effectiveness of the dataset and acceleration strategy in the pipeline through ablation experiments. We also conducted further experiments to test the scalability of the pipeline in the ablation experiments. Additionally, we performed comprehensive comparative experiments to evaluate the pipeline’s performance on downstream tasks.
Pre-training Setup. Our pre-training experiment setup largely follows MAE [10]. The masking ratio is 85%, and the reconstruction ratio is 25%. For patch selection, the adjustment threshold is set at 300 epochs, with the last 200 epochs using random distance to enhance robustness.
4.1 Ablation Study
To validate the pipeline’s effectiveness, we conducted separate evaluations of both the OpticalRS-4M dataset generated by the pipeline and the accelerated training method, SelectiveMAE. Additionally, we assessed the pipeline’s scalability in collecting and processing datasets of varying sizes.
4.1.1 Superiority of OpticalRS-4M
Table 1: Comparison of downstream task performance at matched data throughput between MillionAID and OpticalRS-4M, assessed using MAE [10] pre-training and ViT Base models [39].
Dataset Method Images Number Epoch Sence Object Semantic
Classification Detection Segmentation
AID RESISC-45 DIOR DIOR-R LoveDA SpaceNetv1
OA (TR=20% / 50%) OA (TR=10% / 20%) $\mathrm{mAP}_{50}$ $\mathrm{mAP}_{50}$ mIoU mF1
MillionAID MAE 1 million 800 94.92/97.38 89.20/93.60 71.80 62.33 51.24 79.24
OpticalRS-4M MAE 1 million 800 96.26/97.98 91.53/93.88 75.90 66.46 53.10 79.35
OpticalRS-4M MAE 2 million 400 96.64/98.10 91.80/94.31 73.90 65.95 52.86 79.37
OpticalRS-4M MAE 3 million 267 96.67/98.18 92.24/94.41 75.40 67.07 52.39 79.37
OpticalRS-4M MAE 4 million 200 96.10/98.03 92.38/94.30 74.70 66.26 52.75 79.23
Dataset Diversity. In self-supervised pre-training of RSFMs, the primary goal is to enhance downstream task performance. We define dataset diversity as its ability to improve various types of these tasks. While datasets like MillionAID focus on classification, our OpticalRS-4M dataset is designed to extend the benefit to detection and segmentation tasks. We assess this by evaluating the effectiveness of models trained on different datasets with equivalent data throughput. For instance, MillionAID, with 1 million samples, is trained using the basic MAE method for 800 epochs as a benchmark. To showcase the diversity of OpticalRS-4M, containing 4 million samples, we train it with MAE for just 200 epochs. Additionally, we validate OpticalRS-4M’s diversity by progressively reducing its training data — down to 3 million, 2 million, or even 1 million samples — while keeping data throughput constant. As shown in Table 1, our experiments demonstrate that even when randomly sampling 1 million images from OpticalRS-4M, it outperforms MillionAID in downstream tasks, particularly in object detection and segmentation. This consistent superior performance, even with fewer images, highlights OpticalRS-4M’s greater diversity and effectiveness, emphasizing the benefits of using larger and more varied datasets for unsupervised pre-training.
The results in Table 1 indicate that the optimal outcomes at equivalent data throughput are achieved with 3 million samples rather than 4 million. In most cases, the model pre-trained on 3 million samples outperforms across various downstream tasks. Notably, on the DIOR and LoveDA datasets, the best results were obtained with 1 million samples trained for 800 epochs. However, as shown in Table 2, when we extended the training on 4 million samples to more epochs, the model achieved the highest average performance across most downstream tasks. This suggests that training on 4 million samples for only 200 epochs was insufficient for the MIM method to fully leverage the large dataset, resulting in inferior generalization in downstream tasks.
Table 2: Performance on downstream tasks with extended training on OpticalRS-4M.
Dataset Method Images Number Epoch Sence Object Semantic
Classification Detection Segmentation
AID RESISC-45 DIOR DIOR-R LoveDA SpaceNetv1
OA (TR=20% / 50%) OA (TR=10% / 20%) $\mathrm{mAP}_{50}$ $\mathrm{mAP}_{50}$ mIoU mF1
OpticalRS-4M MAE 3 million 267 96.67/98.18 92.24/94.41 75.40 67.07 52.39 79.37
OpticalRS-4M MAE 4 million 100 95.12/97.54 91.26/93.85 - - - -
OpticalRS-4M MAE 4 million 200 96.10/98.03 92.38/94.30 74.70 66.26 52.75 79.23
OpticalRS-4M MAE 4 million 400 96.36/98.16 92.41/94.03 - - - -
OpticalRS-4M MAE 4 million 800 96.88/98.22 92.44/94.43 75.40 67.35 52.80 79.41
Extended training on OpticalRS-4M. To maximize the potential of OpticalRS-4M, we extended training using the MAE pre-training method and ViT-B, varying only the number of training epochs while keeping other settings consistent. The fine-tuning results, shown in Table 2, reveal that as the number of epochs increases, OpticalRS-4M shows greater advantages in downstream tasks. This indicates that extended training is necessary to fully harness OpticalRS-4M’s capabilities. However, the gains diminish with additional epochs, suggesting that merely increasing epochs may not suffice. To unlock the full potential of OpticalRS-4M, exploring more effective pre-training methods, such as the proposed SelectiveMAE, may be crucial.
Table 3: Comparison of downstream task performance using SelectiveMAE and MAE [10]. The training time is measured using 8 NVIDIA A100 GPUs.
Method Dataset Epoch Total Training Sence Object Semantic
Classification Detection Segmentation
Hours AID RESISC-45 DIOR DIOR-R LoveDA SpaceNetv1
OA (TR=20% / 50%) OA (TR=10% / 20%) $\mathrm{mAP}_{50}$ $\mathrm{mAP}_{50}$ mIoU mF1
MAE OpticalRS-4M 800 177 96.88/98.22 92.44/94.43 75.40 67.35 52.80 79.41
SelectiveMAE OpticalRS-4M 800 86 (2.1×2.1$\times$2.1 ×) 96.78/98.12 93.35/94.58 75.70 67.78 53.05 79.50
4.1.2 Effectiveness of SelectiveMAE
Performance Advantage. To further validate the accelerated pre-training strategy of our pipeline, we conducted extensive experiments on OpticalRS-4M using both MAE [10] and SelectiveMAE. The performance of ViT-B on various downstream tasks after pre-training is presented in Table 3. The results indicate that after pre-training on OpticalRS-4M, SelectiveMAE not only outperforms MAE but also achieves this at 2.1 times the speed. This makes SelectiveMAE particularly well-suited for large-scale datasets, offering significant time savings during training on OpticalRS-4M.
Table 4: The accelerating performance of SelectiveMAE for different backbones compared to MAE [10]. The training time is measured using 8 NVIDIA A100 GPUs.
Model Backbone Params Total Training Hours
MAE ViT-B 86M 101
SelectiveMAE ViT-B 86M 48 (2.1×2.1$\times$2.1 ×)
MAE ViT-L 307M 113
SelectiveMAE ViT-L 307M 50 (2.3×2.3$\times$2.3 ×)
Pre-training Efficiency. To further highlight the efficiency advantage of SelectiveMAE, we evaluate the pretraining time with different backbones on the MillionAID dataset, as shown in Table 4. It can be seen that, since only 40% of the image patches (15% for the encoder and 25% for the decoder) are involved in the calculation, our methods present more than doubled acceleration for both base and large ViT models compared to the vanilla MAE baseline, where the acceleration advantage is more significant for larger sized models.
4.1.3 Scalability of the Pipeline
Table 5: Training cost (measured using 8 NVIDIA A100 GPUs) of SelectiveMAE and MAE [10] on OpticalRS-4M with different data volumes.
Method Images Number Epoch Total Training Hours
MAE 1 million 800 54
SelectiveMAE 1 million 800 25 (-29)
MAE 4 million 800 177
SelectiveMAE 4 million 800 86 (-81)
To evaluate the scalability of our pipeline, we organized datasets containing 1 million and 4 million samples and conducted experiments on each. The results, presented in Table 5, show that with 1 million samples, the proposed SelectiveMAE method reduced training time by 29 hours. When the dataset was expanded to 4 million samples, the time savings increased to 81 hours. These findings demonstrate that our pipeline scales effectively with larger datasets, offering significant pre-training acceleration.
4.2 Main Results
4.2.1 Fine-tuning Performance on Downstream Tasks